OpenStudy (eboogie4):

What is the diameter of a circle with a circumference of 62.83185? A. 5 units B. 10 units C. 20 units D. 25units

pooja195 (pooja195):

Divide 62.83185 by pi

pooja195 (pooja195):

\[\huge~\rm~\bf~\frac{ 62.83185 }{ 3.14 }\]

OpenStudy (jdoe0001):

lol 24 what the wingspanens \(\bf circumference=2\pi r\qquad \qquad r=radius=\frac{diameter}{2}\implies 2r=diameter \\ \quad \\ circumference=\pi 2r\implies circumference=\pi \cdot diameter \\ \quad \\ \cfrac{circumference}{\pi }=diameter\)

OpenStudy (bigdaddyharsha):

C